A study on the opinion of the dicoradores regarding the most desirable primary color for use in executive offices showed the following results:

Primary color

Number of decorators

Red

92

Orange

86

Yellow

46

Green

91

blue

37

Indigo

46

Violet

two


What is the probability that a decorator does not prefer blue?

a. 0.8850

b. 0.9075

c. 0.7725

d. 1.000
Answer #1
Red 92
Orange 86
Yellow 46
Green 91
Blue 37
Indigo 46
Violet 2
Total 400

The total number of decorators are 400.

Blue decorators are 37

P( Prefer blue) = 37/400 = 0.0925

The probability that a decorator does not prefer blue = 1 - P(Prefer blue) = 1 - 0.0925 = 0.9075

b. 0.9075